OpenStudy (anonymous):

how do you sımplıfy sq.rt.1-4sın(x)^2?

OpenStudy (anonymous):

I think you can do something like this. first, I let sqr aside, I will be back to it later. I have 1-4sin^2(x) = 1-2(1-cos(2x) =2cos(2x)-1 .Are you with me then?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and cos2x = 2cos^2(x)+1 replace to 2cos (2x) and, now is your turn, do it.

OpenStudy (anonymous):

I am not sure I understand what you are askıng me to substıtute where?

OpenStudy (anonymous):

4cos^2(x)?

OpenStudy (anonymous):

ok, wait few minutes. I don't do it yet., just give out my idea. but if you stuck, let me try. 50% believe on my stuff only, please.

OpenStudy (anonymous):

where I am I supposed to substıtute what?

OpenStudy (anonymous):

@niksva take it over. don't just send me something in secret. contribute, please

OpenStudy (anonymous):

I don't know why, but I got a negative value of square root. It cannot be, right?

OpenStudy (anonymous):

@jtthoma4 and @Hoa can we replace cos2x = cos^2(x) - sin^2(x) and 1 = sin^2(x) + cos^2(X) in the second step ? hope it will simplify the prob

OpenStudy (anonymous):

@niksva I try, but it goes around and cannot be simplified still.

OpenStudy (anonymous):

did you guys check the domain of the function? \[1-4\sin^2(x)\ge0\]

OpenStudy (anonymous):

ıs thıs a valıd statement; 2sq.rt.1-sın(x)^2

OpenStudy (anonymous):

are you doing some sort of integration or differentiation? and you want to make this a simple one?

OpenStudy (anonymous):

@hoa after replacing i m getting \[\sqrt{\cos ^2(x)- 3\sin ^2(x)}\]

OpenStudy (anonymous):

yeah, me too and I break 3sin^2 = sin^2 +2 sin^2 . and then cos^2 -sin^2 =1. I got 1-2sin^2 . you do next,

OpenStudy (anonymous):

sorry, I am wrong at that stuff, cos^2 -sin^2 = cos 2x

OpenStudy (anonymous):

and 2sin^2(x) = 1-cos 2x . . continue please

OpenStudy (anonymous):

why do you bother? the person who asked the question does not even care!

OpenStudy (anonymous):

@electrokid well said